OpenStudy (anonymous):

Evaluate the limit, if it exists. (If an answer does not exist, enter DNE.) lim h → 0 (49 + h)^.5 − 7) / h

OpenStudy (anonymous):

i get 98+h over h^2

OpenStudy (anonymous):

wow a decimal exponent!!

OpenStudy (anonymous):

\[\frac{\sqrt{49+h}-7}{h}\] is what you really have the gimmick here is to rationalize the numerator by multiplying by \[\frac{\sqrt{49+h}+7}{\sqrt{49+h}+7}\]

OpenStudy (anonymous):

leave the denominator in factored form, i.e. don't multiply out

OpenStudy (anonymous):

wanna try it or you still need help?

OpenStudy (anonymous):

i think I got it. Thanks!

OpenStudy (anonymous):

yw let me know what you get and i will check it btw don't be too confused by these in a week you will be able to do this problem instantly in your head using a shortcut method

OpenStudy (anonymous):

??? whats the method?

OpenStudy (anonymous):

it is a secret i can't tell you

OpenStudy (anonymous):

do i count \[(\sqrt{49+h}-7)\times(\sqrt{49-h}+7 )\] as 3 variables?

OpenStudy (anonymous):

times 3 variables

OpenStudy (anonymous):

ok in a week or two you will recognize this as the "derivative" of \(f(x)=\sqrt{x}\) evaluated at \(x=49\) NO NO NO hold on

OpenStudy (anonymous):

if i wrote \(\sqrt{49-h}\) it was a typo it should be \[\frac{\sqrt{49+h}-7}{h}\times \frac{\sqrt{49+h}+7}{\sqrt{49+h}+7}\]

OpenStudy (anonymous):

the numerator uses \((a-b)(a+b)=a^2-b^2\) which is what makes this trick work

OpenStudy (anonymous):

okay so I get 1 over (49+h)^.5 +7

OpenStudy (anonymous):

you should get \[49+h-49=h\] in the numerator

OpenStudy (anonymous):

limit is 1/14?

OpenStudy (anonymous):

then dividing by \(h\) gives you \[\frac{1}{\sqrt{49+h}+7}\] yes

OpenStudy (anonymous):

i mean "yes \(\frac{1}{14}\) is correct"

OpenStudy (anonymous):

are there online videos that can teach me calculus?

OpenStudy (anonymous):

for free?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

give me a second and i can link to one or two

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

first one is not too dynamic, but decent second one you pick the video and the guy shows real clearly and simply how to solve a problem but like i said, don't get too bogged down with these, you will leave this method behind very rapidly

OpenStudy (anonymous):

\[(\frac{ 1}{ 9 }+\frac{ 1 }{ x })/9+x\]

OpenStudy (anonymous):

limit is dne?

OpenStudy (anonymous):

oh no probably not

OpenStudy (anonymous):

i take it it is \[\lim_{x\to 0}\frac{\frac{1}{9}+\frac{1}{x}}{9+x}\] is that right? or with a minus sign?

OpenStudy (anonymous):

gotta add up in the numerator

OpenStudy (anonymous):

oh I got it!

OpenStudy (anonymous):

1/-81

OpenStudy (anonymous):

you get \[\frac{\frac{x+9}{9x}}{9+x}\]

OpenStudy (anonymous):

i think just \(\frac{1}{81}\) if you do not have a minus sign there

OpenStudy (anonymous):

oh wait, was it the limit as \(x\to -9\) then you are correct